OpenStudy (anonymous):

If you have 5x - 4y = 3 and y = 2x +4, then use substitution.

OpenStudy (anonymous):

montano please help If you have 5x - 4y = 3 and y = 2x +4, then use substitution.

OpenStudy (anonymous):

Okay. So you have 5x - 4(2x + 4) =3 using substitution. This gives an expanded form of: 5x - 8x -16 = 3 Simplified we have: -3x = 13 x = -19/3 If this helps, please leave a medal.